OpenStudy (anonymous):

SOMEBODY ANYBODY! HELP A SENIOR GRADUATE! Larry grows cabbages in a square plot in his garden. He created a drawing of his cabbage plot on a coordinate grid. The locations of three corners of the plot are represented by coordinates (7, 4), (5, 7), and (2, 5). Which of these coordinates represent the fourth corner? (3, 1) (4, 2) (3, 2) (4, 1)

OpenStudy (amistre64):

best bet is to plot it out and get a feel for the last corner

OpenStudy (amistre64):

we could also turn these into vector and make it easier ... maybe

OpenStudy (anonymous):

Thats the thing im really bad at graphs i was hopping someone could just answer it and tell how they did it :(

OpenStudy (anonymous):

what is vector?

OpenStudy (amistre64):

lets avoid the "just answering" scenario and see if we can work our way to a suitable answer

OpenStudy (amistre64):

graphing points shouldnt be that hard, the numbers even yell you where it goes

OpenStudy (anonymous):

ok do you get exactly what there asking?

OpenStudy (amistre64):

i do, we want a square and they give 3 points and want the fourth

OpenStudy (amistre64):

lets find some opposite corners to play with

OpenStudy (anonymous):

ok im going to try

OpenStudy (amistre64):



OpenStudy (amistre64):

lets find the midpoint of these 2 points and that will give us a center to work from

OpenStudy (amistre64):

add the points and divide by 2

OpenStudy (amistre64):

2,5 7,4 ---- 9,9 /2 = 4.5, 4.5 our center point, and midpoint between corners is then 4.5, 4.5

OpenStudy (amistre64):

we can use this with the other point to get to our corner 5 , 7 x, y -------- 4.5 , 4.5 (5+x)/2 = 4.5 5+x = 9 when x=4 7+y = 9 when y=2 i would say that our missing point is that 4,2

OpenStudy (amistre64):

i just realized something after reading this thing 2 months after the fact :) 2,5 5,7 7,4 x,y --- ---- 9,9 9,9 i knew each set of points had to have a common midpoint, so in hindsight, we just needed to determine (x,y) that would add up to the same amounts :) (x,y) -> (4,2)